OpenStudy (anonymous):

given sin u=-5/13,pi

OpenStudy (anonymous):

help plzzz

OpenStudy (anonymous):

\[u =\sin^{-1} (-5/13) = - 22.62 \deg\] But we're given π<u<3π/2. So, u must be π + 22.62 deg or \[u = 180 \deg + 22.62 \deg = 20.2.62 \deg or 3.536 rad\] csc v = sqrt(10/3) or sin v = sqrt(3/10) Now, \[v = \sin^{-1} \sqrt{3/10} = 33.211 \deg or 0.58 rad\] But we're given π/2<v<π. So, v must be π - 33.211 deg \[v = 180 \deg - 33.211 \deg = 146.789 \deg or 2.562 rad\] \[\cos (u-v) = \cos (3.536 - 2.562) = \cos (0.974) = 0.562\] Your answer is 0.562

OpenStudy (anonymous):

the answer is in fractions tho

OpenStudy (anonymous):

Well, 0.562 = 562/1000 = 281/500. Is it equivalent?

OpenStudy (anonymous):

no like-3sqrt10 over 30

OpenStudy (anonymous):

please write that using the equation button

OpenStudy (anonymous):

\[3\sqrt{10}\div30\]